OpenStudy (anonymous):

How would you find the length of the curve x = t^2 y = t^3 bounds are 0 < t < 2.

OpenStudy (anonymous):

The length of the curve is\[\int\limits_{0}^{2}|v(t)|dt\]

OpenStudy (blockcolder):

Since the curve has given parametric equations, the arc length formula becomes: \[\int_{0}^{2} \sqrt{\left (\frac{dx}{dt}\right)^2+\left(\frac{dy}{dt}\right)^2}dt\]

OpenStudy (anonymous):

first evaluate 't' froom the equation, then compare t, so u can find relation berween x and y like x=y^(2/3), then draw a graph, u can find a length of curve,

OpenStudy (anonymous):

\[ \int_0^2 \sqrt{x'(t)^2+y'(t)^2} \, dt=\int_0^2\sqrt{9 t^4+4 t^2}dt=\\ \int_0^2t \sqrt{9 t^2+4}dt=\\ \] It is easy now. Do the integral by substitution, you should get \[\frac{8}{27} \left(10 \sqrt{10}-1\right) \]